# Scattering of the defocusing Calogero–Moser derivative nonlinear Schrödinger equation

### v1

@@ Abstract
Using the Gérard-type explicit formula, we prove the scattering result of solutions to this equation with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. We also characterize the scattering using the distorted Fourier transform associated with the Lax operator. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases.
@@ 1 Introduction
In [19], Kim and Kwon established the soliton resolution results for both finite-time blow-up solutions and global-in-time solutions to (1.3). More precisely, they proved scattering of solutions to (1.3) in $H^{1,1}(\mathbb{R}):=\{u\in H^{1}(\mathbb{R}):xu\in L^{2}(\mathbb{R})\}$ with small mass initial data ($\|u_{0}\|_{L^{2}}^{2}<2\pi$). Following their approach, it is expected that solutions to (1.1) in $H^{1,1}(\mathbb{R})$ also scatter as time tends to infinity.
In this paper, we applied the Gérard-type explicit formula to prove the scattering result for solutions to the defocusing (CM-DNLS) (1.1) with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. Our analysis also relies on the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$, and we can characterize the scattering term using this distorted Fourier transform. One can see also [10, Section 4] for the analogous result on the distorted Fourier transform corresponding to the focusing (CM-DNLS) (1.3).
@@ 1.1 The Lax pair
In this paper, we denote by $L_{+}^{2}(\mathbb{R})$ the Hardy space corresponding to $L^{2}(\mathbb{R})$ functions having a Fourier transform supported in the domain $\xi\geq 0$. Recall that the space $L_{+}^{2}(\mathbb{R})$ identifies to holomorphic functions on the upper-half plane $\mathbb{C}_{+}:=\{z\in\mathbb{C}:\operatorname{Im}(z)>0\}$ such that
@@
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\left|\widehat{f}\left(0^{+}\right)\right|^{2}=-4\pi\operatorname{Im}\left\langle X^{*}f\mid f\right\rangle\leq 4\pi\|X^{*}f\|_{L^{2}}\|f\|_{L^{2}}.$
@@
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\quad I_{+}(f):=\widehat{f}\left(0^{+}\right).$ (1.6)
@@ 1.3 Main result
Let $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{w\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. Then the solution $u$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
@@
As mentioned in the introduction, Kim and Kwon proved the scattering result for solutions to (1.3) in $H^{1,1}(\mathbb{R})$ [19]. More precisely, their appraoch relies on the pseudo-conformal symmetry, and thus requires $xu_{0}\in L^{2}(\mathbb{R})$. Also, it is expected that we can follow their approach to show the scattering result for (1.1) in $H^{1,1}(\mathbb{R})$. The great progress we made in this paper is to extend the scattering result to more singular chiral initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. One can observe that the (unique) ground state solution $\frac{\sqrt{2}}{x+i}$ of the focusing (CM-DNLS) (1.3) (see [11, Lemma 4.1]) is included in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\frac{1}{4}<\alpha<\frac{1}{2}$ but not in $H^{1,1}(\mathbb{R})$. On the other hand, we can characterize the scattering term using the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$ of (1.1). Furthermore, this is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. See also [6] for the study on the long-time behavior of the Benjamin–Ono equation with rational initial data and [12] for the study on the long-time behavior of the half-wave maps equation with rational initial data.
@@
Following the strategy in this paper, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [10], we can also establish the scattering result for solutions to (1.3) with initial data in $H_{+}^{1,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ for $\alpha>\frac{1}{4}$.
@@ 2.1 Generalized eigenfunctions
Let $u\in H_{+}^{1}(\mathbb{R})$, then $\operatorname{ker}\left(L_{u}-\lambda\right)\cap H_{+}^{1}(\mathbb{R})=\{0\}$.
@@
$\mathcal{M}_{u}(\lambda)=\bigl\{\,m\in L^{\infty}_{+}(\mathbb{R}):\;(L_{u}-\lambda)m=0\ \text{in the distributional sense}\bigr\}$
@@ 2.2 The distorted Fourier transform
If moreover $f\in L^{2}(\mathbb{R})$, then $\widetilde{f}\in L^{2}([0,+\infty))$. Denote by $\mathcal{H}_{c}\left(L_{u}\right)$ the continuous space of $L_{u}$, namely the space of those $f\in L_{+}^{2}(\mathbb{R})$ having a spectral measure with no pure point part. Denote by $P_{c}$ the orthogonal projector from $L_{+}^{2}$ to $\mathcal{H}_{c}\left(L_{u}\right)$.
@@
$\int_{\mathbb{R}}\left|f(x)\right|^{2}dx=\int_{\mathbb{R}}\left|P_{c}f(x)\right|^{2}dx=\frac{1}{2\pi}\int_{0}^{\infty}|\widetilde{f}(\lambda)|^{2}d\lambda.$ (2.2)
@@
From the spectral theorem and the fact that $L_{u}$ has no eigenvalue on $0,\infty)$, we have, for every $f\in L_{+}^{2}(\mathbb{R})$,
$\|f\|_{L^{2}}^{2}=\left\|P_{c}f\right\|^{2}=\frac{1}{2i\pi}\lim_{\varepsilon\rightarrow 0^{+}}\int_{0}^{\infty}\left\langle\left(L_{u}-\lambda-i\varepsilon\right)^{-1}f-\left(L_{u}-\lambda+i\varepsilon\right)^{-1}f,f\right\rangle d\lambda.$ (2.3)
@@
1. $(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{\infty}$ is uniformly bounded and strongly convergent to $(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{\infty}$ as $\varepsilon\rightarrow 0^{+}$.
2. $uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{1}$ is uniformly compact as $\varepsilon\rightarrow 0^{+}$, and is convergent to $uT_{\bar{u}}(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{1}$ as $\varepsilon\rightarrow 0^{+}$.
@@
Applying the spectral theorem to the self-adjoint operator $L_{u}$ restricted to $H_{c}(L_{u})$, combined with Proposition 2.2, we can infer the surjectivity of the distorted Fourier transform.
@@
Now we go to analyse the form of $\partial_{\lambda}a$. We know that, the problem $(L_{u}-\lambda)m_{-}=0,\ell_{m_{-}}(-\infty)=1$ is equivalent to
@@
To obtain $c$, we need to give the aysmptotic behavior of $\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$ at $x=$-$\infty$. For this purpose, we establish the following lemma.
@@
Even in the distorted Fourier transform, the resolvent of $X^{*}f+2tL_{u_{0}}$ is still difficult to compute. However, we can easily compute the resolvent of $i\partial_{\lambda}+2t\lambda$ with the following form
@@
For every $z\in\mathbb{C}_{+}$, it is not hard to see that $\widetilde{h}\in H^{1}(0,\infty),h\in H_{+}^{1}(\mathbb{R})$ and $h\in\operatorname{Dom}(X^{*})$ using the Schur test. Now we justify $I_{+}(h)$. From (2.10), we have
@@
where $\widehat{v}=\widetilde{u}$.
From the density of $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$, we can obtain (2.12) for $u\in L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$.
@@ 3 Long-time behavior
Since $z\in\mathbb{C}_{+}$, Now we apply the Van der Corput lemma and we can conclude that
@@
Since $\|u(t)\|_{L^{2}}=\|u_{0}\|_{L^{2}}=\|\mathrm{e}^{it\partial_{x}^{2}}v_{0}\|_{L^{2}}$, we conclude that this weak limit is exactly the strong limit in $L_{+}^{2}(\mathbb{R})$. The proof is complete. ∎

### v2

@@ Abstract
Using the Gérard-type explicit formula, we prove the scattering result of solutions to this equation with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. We also characterize the scattering term using the distorted Fourier transform associated with the Lax operator. Following our approach developed in this paper, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases.
@@ 1 Introduction
In [19], Kim and Kwon established the soliton resolution results for both finite-time blow-up solutions and global-in-time solutions to (1.3). More precisely, they proved the scattering of solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R}):=\{u\in H^{1}(\mathbb{R}):xu\in L^{2}(\mathbb{R})\}$ with small mass initial data ($\|u_{0}\|_{L^{2}}^{2}<2\pi$) and the soliton resolution for global solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R})$. Following their approach, it is natural to expect that solutions to the defocusing (CM-DNLS) (1.1) in $H^{1,1}(\mathbb{R})$ also scatter as time tends to infinity.
In this paper, we applied the Gérard-type explicit formula to prove the scattering result for solutions to the defocusing (CM-DNLS) (1.1) with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. This is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. Our analysis also relies on the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$, and we can characterize the scattering term using this distorted Fourier transform. Following our approach developed in this paper, combined with the distorted Fourier transform corresponding to the focusing (CM-DNLS) (1.3) introduced in [10, Section 4], we can also conclude the asymptotic bound-state/radiation decomposition for global solutions to the focusing (CM-DNLS) (1.3) in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$.
@@ 1.1 The Lax pair
In this paper, we denote by $L_{+}^{2}(\mathbb{R})$ the Hardy space corresponding to $L^{2}(\mathbb{R})$ functions having a Fourier transform supported in the domain $\xi\geq 0$. Also, we denote by $L_{+}^{p}(\mathbb{R})$($1\leq p\leq\infty$) the space corresponding to $L^{p}(\mathbb{R})$ functions having a Fourier transform supported in the domain $\xi\geq 0$. Recall that the space $L_{+}^{2}(\mathbb{R})$ identifies to holomorphic functions on the upper-half plane $\mathbb{C}_{+}:=\{z\in\mathbb{C}:\operatorname{Im}(z)>0\}$ such that
@@
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\left|\widehat{f}\left(0_{+}\right)\right|^{2}=-4\pi\operatorname{Im}\left\langle X^{*}f\mid f\right\rangle\leq 4\pi\|X^{*}f\|_{L^{2}}\|f\|_{L^{2}}.$
@@
$\forall f\in\operatorname{Dom}\left(X^{*}\right),\quad I_{+}(f):=\widehat{f}\left(0_{+}\right).$ (1.6)
@@
1.2 Global well-posedness
We have the following global-well-posedness result for the defocusing (CM-DNLS) (1.1).
Theorem 1.1 ([11, 17]).
Given $u_{0}\in H_{+}^{s}(\mathbb{R})$ with $s\geq 0$, there exists a solution $u\in C(\mathbb{R};H_{+}^{s}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0,x)=u_{0}(x)$.
Remark 1.2.
We have also the analogous global well-posedness result in $H_{+}^{s}(\mathbb{R}))$ ($s\geq 0$) for the focusing (CM-DNLS) (1.3) with small mass initial data ($\|u\|_{L^{2}}^{2}<2\pi$, or $u_{0}\in H_{+}^{1}(\mathbb{R})$ with $\|u_{0}\|_{L^{2}}^{2}=2\pi$) [11, 17]. For the focusing (CM-DNLS) (1.3) with large mass initial data ($\|u\|_{L^{2}}^{2}>2\pi$), the finite-time blow up phenomena would occur [18].
@@ 1.4 Main result
Let $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{w\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}w\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. Then the solution $u\in C(\mathbb{R};H_{+}^{1}(\mathbb{R}))$ to the defocusing (CM-DNLS) (1.1) with $u(0)=u_{0}$ satisfies
@@
As mentioned in the introduction, Kim and Kwon proved the scattering result for solutions to (1.3) in $H^{1,1}(\mathbb{R})$ [19]. More precisely, their appraoch relies on the pseudo-conformal symmetry, and thus requires $xu_{0}\in L^{2}(\mathbb{R})$. Also, it is expected that we can follow their approach to show the scattering result for the defocusing (CM-DNLS) (1.1) in $H^{1,1}(\mathbb{R})$. The great progress we made in this paper is to extend the chiral scattering result to more singular chiral initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ with $\alpha>\frac{1}{4}$. On the other hand, we can characterize the scattering term using the distorted Fourier transform associated with the Lax operator $L_{u_{0}}$ of the defocusing (CM-DNLS) (1.1). In fact, following the method used in this paper, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [10, Section 4], we can also establish the scattering result for solutions to the focusing (CM-DNLS) (1.3) with initial data in $H_{+}^{1,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ for $\alpha>\frac{1}{4}$, which is an extension of the chiral scattering result for the focusing (CM-DNLS) (1.3) in [19]. See Remark 3.2 for details. Another highlight of this paper is that it is one of the first works that apply the Gérard-type explicit formula to study the long-time behavior of an integrable equation for a broad class of initial data, beyond the previously studied rational cases. See also [6] for the study on the long-time behavior of the Benjamin–Ono equation with rational initial data and [12] for the study on the long-time behavior of the half-wave maps equation with rational initial data.
@@
In [19], Kim and Kwon also showed the soliton resolution for global solutions to the focusing (CM-DNLS) (1.3) in $H^{1,1}(\mathbb{R})$. With a slight modification of the proof of Theorem 1.5, combined with the analogous result on the distorted Fourier transform associated with the focusing (CM-DNLS) (1.3) [10, Section 4], we can conclude the asymptotic bound-state/radiation decomposition for global solutions to (1.3) in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$, i.e., if $u\in C(\mathbb{R};H_{+}^{1}(\mathbb{R}))$ is a global solution to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}\in H_{+}^{1,\alpha}(\mathbb{R})$ for $\alpha>\frac{1}{4}$, then
$u(t,x)=\text{bound-states}+\mathrm{e}^{i\partial_{x}^{2}}v_{0}+o(1)\text{ as }t\to\infty$
in $L_{+}^{2}(\mathbb{R})$. Here $v_{0}$ can be defined as in Theorem 1.5, using the distorted Fourier transform introduced in [10, Section 4]. To establish soliton resolution in this setting, one would need to show that each bound state asymptotically converges to a traveling soliton profile. See Remark 3.2 and Remark 3.3 for details.
Remark 1.8.
One can observe that the (unique) ground state solution $\frac{\sqrt{2}}{x+i}$ to the focusing (CM-DNLS) (1.3) (see [11, Lemma 4.1]) is in $H_{+}^{1,\alpha}(\mathbb{R})$ for $\frac{1}{4}<\alpha<\frac{1}{2}$ but not in $H^{1,1}(\mathbb{R})$. This shows that our progress on the long-time analysis of (1.1) and (1.3) is crucial.
@@ 2.1 Generalized eigenfunctions
Let $u\in H_{+}^{1}(\mathbb{R})$, then $L_{u}$ has no eigenvalues.
@@
Remark 2.2.
Using the Weyl’s theorem, we can easily conclude that $\sigma_{ess}(L_{u})=\sigma_{ess}(D)=[0,\infty)$. Since $L_{u}$ has no eigenvalues, we can conclude that $\sigma(L_{u})=\sigma_{ac}(L_{u})\cup\sigma_{sc}(L_{u})=[0,\infty)$.
Remark 2.3.
For the Lax operator of the focusing (CM-DNLS) (1.3), if $\|u\|_{L^{2}}^{2}<2\pi$, we can also infer that this Lax operator has no eigenvalues.
@@
$\mathscr{M}_{u}(\lambda)=\bigl\{\,m\in L^{\infty}_{+}(\mathbb{R}):\;(L_{u}-\lambda)m=0\ \text{in the distributional sense}\bigr\}$
@@ 2.2 The distorted Fourier transform
If moreover $f\in L^{2}(\mathbb{R})$, then $\widetilde{f}\in L^{2}([0,+\infty))$. Denote by $\mathcal{H}_{c}\left(L_{u}\right)$ the continuous space of $L_{u}$, namely the space of those $f\in L_{+}^{2}(\mathbb{R})$ having a spectral measure with no pure point part. Denote by $P_{c}(L_{u})$ the orthogonal projector from $L_{+}^{2}$ to $\mathcal{H}_{c}\left(L_{u}\right)$.
@@
$\int_{\mathbb{R}}\left|f(x)\right|^{2}dx=\int_{\mathbb{R}}\left|P_{c}(L_{u})f(x)\right|^{2}dx=\frac{1}{2\pi}\int_{0}^{\infty}|\widetilde{f}(\lambda)|^{2}d\lambda.$ (2.2)
@@
From the spectral theorem and the fact that $L_{u}$ has no eigenvalue, we have, for every $f\in L_{+}^{2}(\mathbb{R})$,
$\|f\|_{L^{2}}^{2}=\left\|P_{c}(L_{u})f\right\|^{2}=\frac{1}{2i\pi}\lim_{\varepsilon\rightarrow 0_{+}}\int_{0}^{\infty}\left\langle\left(L_{u}-\lambda-i\varepsilon\right)^{-1}f-\left(L_{u}-\lambda+i\varepsilon\right)^{-1}f,f\right\rangle d\lambda.$ (2.3)
@@
1. $(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{\infty}$ is uniformly bounded and strongly convergent to $(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{\infty}$ as $\varepsilon\rightarrow 0_{+}$.
2. $uT_{\bar{u}}(D-\lambda\mp i\varepsilon)^{-1}:L^{1}\rightarrow L^{1}$ is uniformly compact as $\varepsilon\rightarrow 0_{+}$, and is convergent to $uT_{\bar{u}}(D-\lambda\mp i0)^{-1}:L^{1}\rightarrow L^{1}$ as $\varepsilon\rightarrow 0_{+}$.
@@
Applying the spectral theorem to the self-adjoint operator $L_{u}$ restricted to $\mathcal{H}_{c}(L_{u})$, combined with Proposition 2.4, we can infer the surjectivity of the distorted Fourier transform.
@@
Remark 2.12.
From the the spectral theorem and Proposition 2.8, we can also conclude that $\sigma(L_{u})=\sigma_{ac}(L_{u})=[0,\infty)$.
@@
2.3 Computation of the scattering term
We have already completed the distorted Fourier transform theory in the previous part. In this section, using the distorted Fourier transform, we try to find the relation between $X^{*}f$ and $i\partial_{\lambda}\widetilde{f}$, and then we can compute the scattering term. First we go to analyze $\partial_{\lambda}a$. We know that, the problem $(L_{u}-\lambda)m_{-}=0,\ell_{m_{-}}(-\infty)=1$ is equivalent to
@@ 2.3 Computation of the scattering term
To obtain $c$, we need to give the aysmptotic behavior of $\mathrm{e}^{-i\lambda x}n_{1}(x,\lambda)$ as $x\to$-$\infty$. For this purpose, we establish the following lemma.
@@
Even in the distorted Fourier transform, the resolvent of $X^{*}f+2tL_{u_{0}}$ is difficult to compute. However, we can easily compute the resolvent of $i\partial_{\lambda}+2t\lambda$ with the following form
@@
For every $z\in\mathbb{C}_{+}$, it is not hard to see that $\widetilde{h}\in H^{1}(0,\infty),h\in H_{+}^{1}(\mathbb{R})$ and $h\in\operatorname{Dom}(X^{*})$ using the Schur test, so we can apply (2.11) for $h$. Now we justify $I_{+}(h)$. From (2.10), we have
@@
where $\widehat{v}(\xi)=\widetilde{u}(\xi)$ for $\xi>0$ and $\widehat{v}(\xi)$ for $\xi<0$. From the density of $\mathcal{S}_{+}(\mathbb{R})$ in $L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$ and (2.11), we can obtain (2.12) for $u\in L_{+}^{\frac{4}{3}}(\mathbb{R})\cap H_{+}^{1}(\mathbb{R})$.
In fact, we can easily see that the scattering term in Theorem 1.5 is exacly $\mathrm{e}^{it\partial_{x}^{2}}v$ obtained in (2.12) with $u=u_{0}$.
@@ 3 Long-time behavior
Since $z\in\mathbb{C}_{+}$, Now we apply the Riemann–Lebesgue lemma and we can conclude that
@@
Since $\|u(t)\|_{L^{2}}=\|u_{0}\|_{L^{2}}=\|\mathrm{e}^{it\partial_{x}^{2}}v_{0}\|_{L^{2}}$, we conclude that this weak convergence is exactly the strong convergence in $L_{+}^{2}(\mathbb{R})$. The proof is complete. ∎
Remark 3.2.
Following the approach developed in Section 2.3 and Section 3, we can also conclude the asymptotic bound-state/radiation decomposition for global solutions with initial data in $H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$.
We fix $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R}):=\{u\in H_{+}^{1}(\mathbb{R}):|x|^{\alpha}u\in L^{2}(\mathbb{R})\}$ for $\alpha>\frac{1}{4}$. From [11, Section 5], we know that every eigenvalue of $L_{u_{0}}=D-u_{0}T_{\bar{u}_{0}}$ (the Lax operator of (1.3)) is simple, and the number $N$ of eigenvalues of $L_{u_{0}}$ is finite with $N\leq\frac{\|u_{0}\|_{L^{2}}^{2}}{2\pi}$. From [10, Theorem 4.3], we know that for every $f\in L_{+}^{2}(\mathbb{R})$
$f(x)=\sum_{j=1}^{N}\left\langle f,\varphi_{j}\right\rangle\varphi_{j}(x)+\mathcal{F}_{d}^{-1}(\widetilde{f})(x)=P_{p}(L_{u_{0}})f(x)+P_{ac}(L_{u_{0}})f(x).$ (3.4)
Here $\left\{\varphi_{j}\right\}_{j=1}^{N}$ are the normalized eigenfunctions of $L_{u_{0}}$, $\mathcal{F}_{d}(f)=\widetilde{f}$ is the distorted Fourier transform defined in [10, (4.1)], $P_{p}(L_{u_{0}})$ denote the projection onto the eigensubspace of $L_{u_{0}}$ and $P_{ac}(L_{u_{0}})$ denote the projection onto the absolutely continuous subspace of $L_{u_{0}}$. Now we apply (3.4) for $f=u_{0}$, and we obtain
$u_{0}(x)=P_{p}(L_{u_{0}})u_{0}(x)+P_{ac}(L_{u_{0}})u_{0}(x).$ (3.5)
Now we assume that the solution $u(t,\cdot)\in H_{+}^{1}(\mathbb{R})$ to the focusing (CM-DNLS) (1.3) with $u(0)=u_{0}$ exists globally in time (this is always true if $\|u_{0}\|_{L^{2}}^{2}\leq 2\pi$). From the explicit formula (3.3), we infer that for every $t\in\mathbb{R}$ and every $z\in\mathbb{C}_{+}$,
$u(t,z)=\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}\varphi\right]+\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}P_{ac}(L_{u_{0}})u_{0}\right].$ (3.6)
For the second term above, following the approach in the proof of Theorem 3.1, we can show that
$\forall z\in\mathbb{C}_{+},\qquad\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}P_{ac}(L_{u_{0}})u_{0}\right]\rightarrow\mathrm{e}^{it\partial_{x}^{2}}v_{0}(z),$
where $\widehat{v_{0}}(\xi)=\widetilde{u_{0}}(\xi)$ for every $\xi>0$ with $\widehat{v_{0}}(\xi)=0$ for every $\xi<0$. From this we can deduce the sacttering term of the focusing (CM-DNLS) (1.3). Notice that if $\|u_{0}\|_{L^{2}}^{2}<2\pi$, $L_{u_{0}}$ has no eigenvalues. In this case $P_{p}(L_{u_{0}})=0$ and $P_{ac}(L_{u_{0}})u_{0}=u_{0}$, so from the above argument we can directly deduce the scattering of the focusing (CM-DNLS) (1.3) for $u_{0}\in H_{+}^{1,\alpha}(\mathbb{R})\cap\{\|u\|_{L^{2}}^{2}<2\pi\}$ with $\alpha>\frac{1}{4}$.
Now we analyse the first term on the right hand side in (3.6). Following the method used in the proof of [8, Lemma 3.1], we can also deduce that
$\frac{1}{2i\pi}I_{+}\left[\left(X^{*}+2tL_{u_{0}}-z\mathrm{Id}\right)^{-1}P_{p}(L_{u_{0}})u_{0}\right]=U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}P_{p}(L_{u_{0}})u_{0}(z).$
and
$u(t)=U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}u_{0}.$
Since $\left\{\varphi_{j}\right\}_{j=1}^{N}$ are the normalized eigenfunctions of $L_{u_{0}}$, we know that
$U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}P_{p}(L_{u_{0}})u_{0}$
$=$ $\,\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}\varphi_{j}$
$=$ $\,\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\mathrm{e}^{-it\lambda_{j}^{2}}U(t)\varphi_{j}.$
Here the unitary operator $U(t)$ is defined by the linear initial value problem in $\mathscr{L}\left(L_{+}^{2}(\mathbb{R})\right)$,
$U^{\prime}(t)=B_{u(t)}U(t),\qquad U(0)=\operatorname{Id}\text{ with }B_{u}:=uT_{\partial_{x}\bar{u}}-\partial_{x}uT_{\bar{u}}+i(uT_{\bar{u}})^{2}.$
From the Lax identity $L_{u(t)}U(t)=U(t)L_{u_{0}}$, we can infer that $\phi_{j}(t):=\mathrm{e}^{-it\lambda_{j}^{2}}U(t)\varphi_{j}$ is the normalized eigenfunctions of $L_{u(t)}$. Thus we have
$U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}P_{p}(L_{u_{0}})u_{0}=\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\phi_{j}(t).$
So we have
$u(t)=\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\phi_{j}(t)+\mathrm{e}^{it\partial_{x}^{2}}v_{0}+o(1)$ (3.7)
in $L_{+}^{2}(\mathbb{R})$. Notice that (3.7) is exactly the asymptotic bound-state/radiation decomposition for the global solution $u(t)$ to the focusing (CM-DNLS) (1.3). In fact, we have
$\left\langle u_{0},\varphi_{j}\right\rangle=\left\langle U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}u_{0},U(t)\mathrm{e}^{-itL_{u_{0}}^{2}}\varphi_{j}\right\rangle=\left\langle u(t),\phi_{j}(t)\right\rangle,$
therefore $\sum_{j=1}^{N}\left\langle u_{0},\varphi_{j}\right\rangle\phi_{j}(t)=P_{p}(L_{u(t)})(u(t))$. Moreover, from [10, Lemma 3.7], we know that for every $1\leq j\leq N$,
$|\left\langle u_{0},\varphi_{j}\right\rangle|^{2}=2\pi.$
Then we can deduce that
$\|v_{0}\|_{L^{2}}^{2}=\|P_{ac}(L_{u_{0}})u_{0}\|_{L^{2}}^{2}=\|u_{0}\|_{L^{2}}^{2}-\sum_{j=1}^{N}|\left\langle u_{0},\varphi_{j}\right\rangle|^{2}=\|u_{0}\|_{L^{2}}^{2}-2\pi N,$
which coincides with one of the statements in [19, Theorem 1.1].
Remark $3.$
From the soliton resolution result in [19, Theorem 1.1] and Remark 3.2, we are naturally led to the following conjecture: as $t\to\infty$, $\sqrt{2\pi}\phi_{j}(t)$ should asymptotically converge to a traveling soliton profile.